### v1

# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using AMPT

Abstract
The study of symmetry plane correlations (SPCs) can be useful in characterizing the direction of the anisotropic emission of produced particles in the final state. The study of SPCs provides an independent method to understand the transport properties of the system formed in heavy-ion collisions. Similar to anisotropic flow coefficients, which are largely influenced by the initial spatial anisotropy, SPCs also depend upon the participant plane correlation measured using the participating nucleons of the collision overlap region. In this paper, SPCs have been studied in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using the event generator AMPT. In addition to their behaviour with the changing centrality of the collision, their event shape dependence has also been studied for the first time, using the event shape classifier transverse spherocity. The Gaussian estimator has been used to evaluate the correlations, and these have been compared to the participant plane correlations defined in an analogous way to the symmetry plane correlations and a qualitative match has been found between them. These event-shape differentiated symmetry plane correlations can be used to deduce the presence of higher-order anisotropies in the initial energy distribution, thus giving insight into the initial geometry of the colliding system, among other applications like model development and model testing using Bayesian analyses.
I Introduction
In the heavy-ion collisions at ultra-relativistic speeds, the formation of a deconfined state of partons (quarks and gluons), called the quark-gluon plasma (QGP), is anticipated [1, 2, 3]. QGP is transient in nature, having a lifetime of the order of $10^{-23}$ seconds [4]. Thus, presently, QGP formation can only be studied through the study of various indirect signatures such as collectivity [6, 5], jet quenching [7], charmonia suppression [8], strangeness enhancement [9] to name a few. QGP is theorized to behave like a perfect fluid. Studies based on anisotropic flow at the RHIC and LHC energies suggest that the ratio of specific shear viscosity to entropy density ($\eta_{s}/s$) is close to $1/4\pi$ which is the lower bound of $\eta_{s}/s$ imposed by AdS/CFT calculations [10, 11, 12, 13, 14, 15, 16]. Further, for a smaller value of $\eta_{s}/s$, the hydrodynamic nature of the QGP can transform the initial state spatial anisotropy of the collision overlap region defined by the participant nucleons into the final state momentum space azimuthal anisotropy of the produced charged particles [17, 5, 18]. The final state azimuthal anisotropy of the charged particles can be quantified using the coefficients of the Fourier expansion of azimuthal distributions of the particles, as follows [18].
$f(\phi)=\frac{1}{2\pi}\left[1+2\sum_{n=1}^{\infty}v_{n}\cos[n(\phi-\psi_{n})]\right]$ (1)
Here, $\phi$ denotes the azimuthal angle in the momentum space. The Fourier coefficients, $v_{n}$, are referred to as the anisotropic flow coefficients, and the parameters $\psi_{n}$ are called the symmetry plane angles. Together, these observables give us an idea of how the particles have been ejected. For example, the second order flow coefficient $v_{2}$ is called the elliptic flow [19] and quantifies the extent up to which the charged particle distribution in the ($p_{x}-p_{y}$) axis can be described by an ellipse. The anisotropic flow coefficients can serve as an instrument to probe for the initial state geometry of the colliding nuclei due to the collective hydrodynamic expansion of the QGP medium [20]. Further discussions regarding the mathematical properties of the flow decomposition of the transverse momentum distribution of the final state particles can be found in Ref. [21].
Some trivial properties of the anisotropic flow coefficients and the symmetry plane angles corresponding to a particular harmonic can be directly inferred from Eq. (1) using the underlying fact that if a change of azimuthal coordinates is made from $\phi\to-\phi$, then the distribution $f(\phi)$, in Eq. (1) remains invariant. Due to this fact, if the Fourier sum in Eq. (1) was to be re-expressed in terms of complex exponential terms, then $v_{n}=v_{-n}$ and $\psi_{n}=\psi_{-n}$. A study of the symmetry plane angles by itself is not feasible as a function of centrality or other event classifiers. This is because, in each collision, there is a random orientation of the reaction plane [22], hence changing the symmetry planes randomly as well. This leads to an event average, $\langle\psi_{n}\rangle=0$. Thus, instead of explicitly studying $\psi_{n}$, it is reasonable to study the correlation between the symmetry plane angles with different event classifiers, including collision centrality. Similar to anisotropic flow coefficients, the correlations of the symmetry plane angles of different orders are sensitive to both initial state correlations and the transport properties of the system. Thus, a study of symmetry plane angles can provide an independent method to understand the initial state transport properties of the medium formed in heavy-ion collisions [23]. In this direction, using the definition of the Gaussian estimator for the symmetry plane correlations (SPC), the correlation among the symmetry planes of different orders has been studied. The computation of symmetry plane angles and of anisotropic flow coefficients involving multi-particle correlations is given below[21]:
$v_{n_{1}}^{a_{1}}v_{n_{2}}^{a_{2}}...v_{n_{k}}^{a_{k}}e^{i(a_{1}n_{1}\psi_{n_{1}}+a_{2}n_{2}\psi_{n_{2}}+...+a_{k}n_{k}\psi_{n_{k}})}$
$=\langle e^{i(n_{1}\phi_{1}+n_{2}\phi_{2}+...+n_{l}\phi_{l})}\rangle$ (2)
In the above equation, the LHS denotes the combination of anisotropic flow coefficients ($v_{n}$) and symmetry plane angles ($\psi_{n}$). The values of the harmonics $n_{i}$ and indices $a_{i}$ are such that $a_{i}$ are positive and hence denote the number of times that a flow term of order $n_{i}$, or equivalently $-n_{i}$ occurs. There is kept independence in the choice of harmonics, which is necessary to define different kinds of correlations appropriately [24], such that the effect of random orientation of the reaction plane from event to event, and the effect of trivial periodicity, i.e., Eq. (1) being invariant under the transformation $\psi_{n}\to\psi_{n}+\frac{2m\pi}{n}$, for integers $0\leq m<n$ play no role when an event average of the symmetry plane correlations is computed. The coefficients are to be chosen so that:
$\sum_{i=1}^{k}a_{i}n_{i}=0$ (3)
The above choice of coefficients ensures that the two caveats necessary to construct a meaningful physical observable to estimate the correlation between different symmetry planes are taken care of. The integer $l$ in the RHS of Eq. (I) denotes the number of particles that need to be correlated in order to compute the LHS. The particle’s final state azimuthal angles $(\phi_{1},\phi_{2},...,\phi_{l})$ are such that all possible multiplets of particles, with appropriate cuts (Section II.3) are considered and their associated harmonics $(n_{1},n_{2},...,n_{l})$ are the same as harmonics $n_{i}$ which appear in the LHS of Eq. (I), $a_{i}$ times. That is, the correlation in the RHS of Eq. (I) is computed using $a_{i}$ different particles associated with the harmonics $n_{i}$.
The collective behaviour, which occurs due to hydrodynamic expansion of the QGP medium, is expected to be sensitive to the initial conditions of the system of the two colliding nuclei [25]. Hence, the various flow coefficients and symmetry planes could give access to the initial configuration of the nuclei if the system has a linear response to its initial conditions. More explicitly, in such a scenario, the SPCs should exhibit similar behaviour to the initial state, participant plane correlations (PPCs), with changing centrality. Further, due to event-by-event fluctuation in the geometry of the collision overlap region, both initial eccentricity and participant plane angle can vary event-by-event. Thus, in a particular centrality class, both anisotropic flow coefficients and symmetry plane angles can have a broad range distribution. This can lead to changes in symmetry plane correlations within a centrality bin. These varying magnitudes of symmetry plane correlations can be studied using event-shape observables. Transverse spherocity [26, 27, 28] is one such event shape observable which can identify the isotropic events, dominated with soft interactions from the jetty events dominated with hard interactions. Studies based on a multi-phase transport model show the applicability of transverse spherocity to separate events in heavy-ion collisions at the LHC energies [32, 31, 30, 29, 20]. The necessity of studying SPCs by classifying collision events by their shapes is obvious since it could give us insight into the variation of the SPCs with the processes that lead to final particle production with a certain geometry. Also, some correlations, for example, which have been experimentally found to be zero, e.g., the correlation of $\psi_{2}$ and $\psi_{3}$ [23], can be verified if they are always zero or happen to be so due to the averaging over events of all shapes, and actually have non-zero correlation in specific event shapes. Such a study will give insight into the usability of different event shapes to probe into correlations amongst symmetry planes. The event shape classifier used for this purpose is transverse spherocity (hereafter referred to as spherocity or $S_{0}$ for brevity).
The paper is organized as follows. We start with a brief introduction in Section I, followed by the description of the methodology of event generation, i.e., the details of the AMPT parameters and the phase space cuts on particle selection for SPC evaluation in Section II. The results are discussed in Section III. Finally, the paper is summarised in Section IV.
II Event Generation and Methodology
In this section, we briefly discuss the event generation using AMPT, transverse spherocity, event and track selection cuts, and the method to calculate the symmetry plane correlations.
II.1 A Multi-phase Transport model (AMPT)
A multi-phase transport model (AMPT) [33] is a Monte Carlo (MC) based transport model that simulates heavy-ion collisions by explicitly dealing with non-equilibrium dynamics of the collision system. AMPT involves four major stages. In the string melting version of AMPT (AMPT-SM), the collision between two nuclei is initialized using HIJING [34, 35], followed by partonic scatterings calculated using Zhang’s Parton Cascade (ZPC) [36], followed by hadronization using quark coalescence model [37], and finally evaluating the hadronic rescatterings using a relativistic transport (ART) framework [38, 39]. In this study, we use AMPT to simulate Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5020$ GeV. The parameters in the Lund symmetric splitting function are set to $a=0.30$ and $b=0.15$ [40]. The parton screening mass is set to 1.2408 $\rm fm^{-1}$ [36]. Flags for shadowing, initial and final state radiations, and random orientation of the reaction plane angle were kept ‘on’. A total of 110,000 minimum bias Pb-Pb collision events are generated for this study using the settings mentioned above.
II.2 Transverse spherocity
Transverse spherocity, which is used for event shape classification, is defined as follows [20, 41, 42]:
$S_{0}=\frac{\pi^{2}}{4}\min_{\hat{n}}\left[\left(\frac{\Sigma_{i}|\vec{p}_{T_{i}}\times\hat{n}|}{\Sigma_{i}|\vec{p}_{T_{i}}|}\right)^{2}\right]$ (4)
Where $\vec{p}_{T_{i}}$ refers to the transverse momentum vector of the charged hadron track $i$ in an event and $\hat{n}$ is a parameter unit vector in the plane transverse to the beam axis which is used to minimize the quantity within the square brackets, and the constant factor at the front is for the purpose of normalization. In accordance with its definition in Ref. [20], only those charged hadronic tracks with transverse momenta $p_{\rm T}>0.15$ GeV/c have been used for the evaluation of spherocity. An additional constraint that has been used is that only those events have been considered for this study, in which there are at least five charged hadrons satisfying the given transverse momentum cut. The extreme limits of $S_{0}\to 0$ and $S_{0}\to 1$ correspond to the event shape being pencil-like and isotropic, respectively. This is mainly because, in a pencil-like event, the particles emerge collimated along a particular direction, and hence, aligning $\hat{n}$ in that direction, all the terms in the numerator of Eq. (4) vanish, leading to $S_{0}\rightarrow 0$. Using the reverse arguments, we have that for an isotropic event, $S_{0}\rightarrow 1$. However, for all events, their $S_{0}$ values lie in between these extreme limits. Further details regarding spherocity can be found in Ref. [20].
II.3 Event and track selection
The evaluation of SPCs and PPCs as defined in Eq. (II.4) and Eq. (II.4) respectively involves in the RHS, averaging the quantities in the angular brackets over different events. Now, each event is classified as per its centrality, which is done using its impact parameter that goes into the input during event generation, using the Glauber Monte Carlo scheme [43], which is used by the HIJING algorithm during nuclear configuration initialization. Further to this centrality classification, each event is subdivided into an event shape class based on the value of its spherocity. There are three event shape classes in each centrality class – low-$S_{0}$ (pencil-like) events (20% events with the least spherocity), high-$S_{0}$ (isotropic) events (20% events with the highest spherocity) and a pool of events with all possible spherocity values, known as $S_{0}$-int events. The event shape and centrality classifications are listed in Table 1.
Centrality [%] b (fm) Low-$S_{0}$ High-$S_{0}$
0-5 0-3.49 0-0.897 0.958-1
5-10 3.49-4.93 0-0.859 0.942-1
10-20 4.93-6.98 0-0.810 0.912-1
20-30 6.98-8.55 0-0.764 0.881-1
30-40 8.55-9.87 0-0.734 0.869-1
40-50 9.87-11 0-0.719 0.869-1
50-60 11-12.1 0-0.718 0.876-1
Table 1: Centrality and spherocity classification of minimum bias Pb-Pb events with $\sqrt{s_{\rm NN}}=5.02$ TeV generated using AMPT. ‘b’ denotes the impact parameter.
After the events have been classified as per their centrality and spherocity classes, the averages in the numerator and denominator in the RHS of Eq. (II.4) are computed using the events that belong to the respective centrality and event shape classes.
The terms inside the angular brackets of the numerator and denominator in the RHS of Eq. (II.4) are obtained for each event. These, in turn, are evaluated by averaging over charged hadron track multiplets as per Eq. (I). For the purpose of averaging over these tracks, only those charged hadronic tracks in the pseudorapidity range $|\eta|<0.8$ and with transverse momenta $5.00.2<p_{\rm T}<$ 5.0 GeV/c are used. These cuts are in order to be consistent with experiment [23] - lower limit owing to ALICE detector efficiency, an upper limit to remove the contribution of non-flow effects that result from jets (which have been taken to be having $p_{\rm T}>5.0$ GeV/c).
II.4 Participant and symmetry plane correlations
From Eq. (I), we can reckon, for a single event, $v_{n}e^{in\psi_{n}}=\langle e^{in\phi}\rangle$. Now, as Pb-Pb collisions at the LHC are associated with high multiplicity, therefore, from event to event, $Re(v_{n})$ and $Im(v_{n})$ (where $Re(z),~{}Im(z)$ denote the real and imaginary part of a complex number $z$ respectively) are expected to have a double Gaussian distribution in the Argand plane [44]. Using this assumption, the correlation between symmetry planes, which are independent of the effect of flow coefficients, has been derived in Ref. [44]. The Gaussian estimator for the symmetry plane correlations is given by [23]:
$\langle\cos(a_{1}n_{1}\psi_{1}+a_{2}n_{2}\psi_{2}+...+a_{k}n_{k}\psi_{k})\rangle_{\rm GE}$
$=\sqrt{\frac{\pi}{4}}\frac{\langle{v_{n_{1}}^{a_{1}}v_{n_{2}}^{a_{2}}...v_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\psi_{1}+a_{2}n_{2}\psi_{2}+...+a_{k}n_{k}\psi_{k})}\rangle}{\sqrt{\left\langle{v_{n_{1}}^{2a_{1}}v_{n_{2}}^{2a_{2}}...v_{n_{k}}^{2a_{k}}}\right\rangle}}$ (5)
Here, the angular brackets denote event averaging over a suitable set of events in a particular event class 1. Further, following its definition in [44], and as explained in Eq. (3), the choice of harmonics and their powers is made such that $\sum_{i=1}^{k}a_{i}n_{i}=0$. An example of its use and evaluation using multi-particle correlations can be found in Ref. [23].
For the sake of comparison to initial state geometry effects, the following definition of eccentricities and participant planes has been used, motivated from its original definition in Ref. [21] and subsequent constraints as established in Ref. [44]:
$\epsilon_{n_{1}}^{a_{1}}\epsilon_{n_{2}}^{a_{2}}...\epsilon_{n_{k}}^{a_{k}}e^{i(a_{1}n_{1}\Phi_{n_{1}}+a_{2}n_{2}\Phi_{n_{2}}+...+a_{k}n_{k}\Phi_{n_{k}})}$
$=\frac{\langle r_{1}^{|n_{1}|}r_{2}^{|n_{2}|}...r_{l}^{|n_{l}|}e^{i(n_{1}\tilde{\phi}_{n_{1}}+n_{2}\tilde{\phi}_{n_{2}}+...+n_{l}\tilde{\phi}_{n_{l}})}\rangle}{\langle r_{1}^{|n_{1}|}r_{2}^{|n_{2}|}...r_{l}^{|n_{l}|}\rangle}$ (6)
In the LHS of this equation, $\epsilon_{n}$ are eccentricities of the initial energy distribution, and $\Phi_{n}$, their corresponding symmetry angles, are called the participant plane angle. The RHS shows the way to compute the combination of $\epsilon_{n}$’s and $\Phi_{n}$’s using $\sum_{i=1}^{k}a_{i}=l$ particles. In the RHS, the angles, $\tilde{\phi}_{i}$’s are the azimuthal angles in the coordinate space, of wounded nucleons. Correspondingly, $r_{i}$, which appear to the power $n_{i}$ in the weight factors, are the radial coordinates in the plane azimuthal to the beam axis of the wounded nucleon whose azimuthal angle $\tilde{\phi}_{i}$ is considered. It is important to note that Eq. (II.4) does not entail the correct definition of $\epsilon_{1}e^{i\Phi_{1}}$[45], as well as it fails to account for the minus signs when it comes to computing just one harmonic order of initial state anisotropies. The absence of these features, however, is not an issue for the study of participant plane correlations, as first-order harmonic has not been considered for participant correlations in this paper, and in addition to that, the choice of harmonics $n_{i}$ would lead to all the negative signs canceling when we account for the rotational invariance of the PPCs under the random orientation of the reaction plane from event to event. In case one wishes to study the participant plane correlations involving the first harmonic, i.e., $n=1$, an explicit $r^{3}$ weight factor must be used [45]. Finally, in analogy to Eq. (II.4), we have the following definition of the Gaussian estimator of the PPCs:
$\langle\cos(a_{1}n_{1}\Phi_{1}+a_{2}n_{2}\Phi_{2}+...+a_{k}n_{k}\Phi_{k})\rangle_{\rm GE}$
$=\sqrt{\frac{\pi}{4}}\frac{\langle{\epsilon_{n_{1}}^{a_{1}}\epsilon_{n_{2}}^{a_{2}}...\epsilon_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\Phi_{1}+a_{2}n_{2}\Phi_{2}+...+a_{k}n_{k}\Phi_{k})}\rangle}{\sqrt{\left\langle{{\epsilon_{n_{1}}^{2a_{1}}\epsilon_{n_{2}}^{2a_{2}}...\epsilon_{n_{k}}^{2a_{k}}}}\right\rangle}}$ (7)
Here, the symbols have their meanings as described previously, and the angular brackets denote event averaging in the respective pool of events, considering their centrality and $S_{0}$ classification.
III Results and Discussion
Firstly, it must be noted that, in systems with high multiplicity, the anisotropic flow coefficients provide the extent to which the distribution of particles in the transverse momentum space looks identical to a regular polygon with $n$ sides, where $n$ is the harmonic order of the flow term being considered. The symmetry plane angle denotes the orientation with respect to a reference, which characterizes the particular harmonic order of anisotropy. This is why the correlation between symmetry plane angles of different harmonic orders of anisotropy will provide independent information that can be obtained from the correlation between anisotropic flow coefficients of the same harmonic orders. One particular reason for this is that the anisotropic flow coefficients typically become smaller with the increase in the order of anisotropy, whereas the symmetry planes need not have any such constraints. Hence, a non-zero value of the SPCs would mean that the symmetry plane angles being considered are correlated among themselves. To understand this with an example, we can consider the combination $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$. Consider that the symmetry planes of harmonic orders $n=6$ and $n=3$ are oriented at angles $\delta_{6}$ and $\delta_{3}$, respectively. Clearly, all possible symmetry planes of order 6 are $\{\delta_{6}+\frac{2k\pi}{6}\}_{k=0}^{5}$. Thus, $6\delta_{6}\in\{6\delta_{6}+2k\pi\}_{k=0}^{5}$. Similarly, $6\delta_{3}\in\{6\delta_{3}+4k\pi\}_{k=0}^{2}$. Hence, $\cos(6(\psi_{6}-\psi_{3}))\in\{\cos(6(\delta_{6}-\delta_{3})+(m+2n)2\pi)\}$ for integers $m,n$. The translation of the argument of cosine function by an integer multiple of $2\pi$ leaves the function invariant, and ultimately we can infer that this correlator shows the correlation between any of the orientations of the $6^{\rm th}$ and $3^{\rm rd}$ order anisotropies in the distribution of the transverse momenta of the final state particles that emerge as the result of a collision. Now, further generalizing, a non-zero value of the GE of the SPCs for a linear combination of symmetry planes $\sum_{i=1}^{k}a_{i}n_{i}\psi_{n_{i}}$, indicates that despite event-by-event fluctuations, the orientations of the harmonics $n_{i}$ to their respective powers of $a_{i}$ are correlated. While for simple choice of harmonics, we can provide direct interpretations in terms of geometry, non-trivial behaviour of correlators with centrality has been observed for different choices of $a_{i}$ while keeping $n_{i}$ same under the constraint that $\sum_{i=1}^{k}a_{i}n_{i}=0$.
Figure 1: Comparison of the centrality dependence of SPCs calculated using various models, including AMPT from the present study, with ALICE measurements in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [46].
III.1 Centrality dependence of SPCs
In this section, before explicitly discussing the SPCs, we begin with the comparison of SPCs calculated using AMPT with experimental measurements. Figure 1 shows the comparison of ALICE measurements with corresponding model calculations. Here, one finds that AMPT provides a good qualitative agreement with experimental results while underestimating the values.
Figure 2: Centrality dependence of symmetry plane correlations of different orders in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT.
Figure 2 shows the symmetry plane correlations of different harmonics as a function of the centrality of collisions in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. In the left panel, which shows the correlations for two symmetry planes, one finds the strength of the correlation is strongest for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$. The strength of the correlation becomes weaker for $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$. Further, it becomes the weakest for $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which is consistent with zero within uncertainties. This decreasing strength of correlation between two symmetry planes is consistent with experimental observations made in Refs. [23, 46]. Interestingly, for collision centrality (0-20)%, one finds that $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$. After, (20-30)centrality, $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ becomes comparable and finally $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ leads the values towards (50-60)centrality. Although, different SPCs show different centrality dependence, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ shows a non-linear centrality dependence. Further, the observed zero value for $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$ indicates zero correlation between $\psi_{2}$ and $\psi_{3}$, although $v_{2}$ and $v_{3}$ show anti-correlation [23, 46, 47, 48, 49, 50, 51, 52, 53, 20]. This hints towards the applicability of SPCs to independently probe the formation of a QCD medium in heavy-ion collisions. Similar to SPCs involving two symmetry planes, one finds different magnitudes for SPCs of three different symmetry planes. Here, $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ shows the strongest correlation and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ has the weakest correlation. Interestingly, $\langle\cos(2\psi_{2}+3\psi_{3}-6\psi_{6})\rangle_{\rm GE}$ has a negative sign, unlike other correlators which are all positive or consistent with zero at the very least. Another interesting observation is that the magnitudes of the SPCs are approximately dependent to the number of particle correlations in the final state. For example, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ each require only three-particle correlations and possesses the largest values among the shown SPCs. In contrast, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ require five and six correlators, respectively, and thus possess a negligibly small value of correlation [23, 46]. One possible explanation is as follows. The fluctuation of flow vectors encoded in the SPCs can be traced back to the initial state fluctuations. The initial state fluctuations can be attributed to the fluctuations due to the finite amount of participants in the collision overlap region. According to the central limit theorem, the average of the random samples converges to a Gaussian distribution with an increase in the number of sampling. For a Gaussian distribution, only the second-order cumulant, namely, the width of the distribution, is non-vanishing. The skewness and kurtosis are small for a distribution close to Gaussian. Here, the order of particle correlations refers to the order of cumulants. Thus, the ordering in magnitudes of SPCs shown in Fig. 2 indicates that the contribution of higher order cumulant (higher order particle correlator) is smaller. On the other hand, the lowest-order cumulants have the leading values. This explains the crossing behaviour between $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$, after (20-30)centrality class, where the number of participating nucleons is smaller. We now move to discuss the geometrical interpretation of each SPC and corresponding PPC.
Figure 3: Centrality dependence of SPCs (upper panel) among $n=2,3,6$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
III.2 Geometrical interpretation of SPCs
III.2.1 SPCs of harmonic orders $n=2,3,6$
The symmetry planes corresponding to the harmonics $n=2,3,6$ represent elliptical, triangular, and hexagonal symmetries, respectively. The correlations amongst them have been shown in Figs. 3(a)-(c). Whenever there is a hexagonal symmetry present in the system, we can deduce trivially from the geometry that choosing any two diametrically opposite vertices of a regular hexagon (that is, those vertices that are separated by the largest distance) would also correspond to a local elliptical symmetry about them. Hence, it can be expected that there is a second-order symmetry plane whose alignment is correlated to the sixth-order symmetry plane. This will be an enhanced effect when $v_{6}$ is more dominant than the second-order anisotropy. In such cases, the symmetry planes $\psi_{6}$ and $\psi_{2}$ are more likely to be correlated. Similarly, by choosing any of the three vertices of a regular hexagon that alternate with each other on the hexagon, we obtain an equilateral triangle. So, a plane about which the transverse distribution has a hexagonal symmetry is also going to be correlated to a plane corresponding to triangular symmetry. Thus, in addition to a strong correlation in $\psi_{6}$ and $\psi_{2}$, a strong correlation between $\psi_{6}$ and $\psi_{3}$ can also be expected for events with higher values of $v_{6}$ than that of $v_{3}$. Again, along the same line of deduction, a higher $v_{6}$ than both $v_{2}$ and $v_{3}$ would also correspond to a higher correlation between $\psi_{2}$ and $\psi_{3}$. Another interesting feature that follows from the presence of hexagonal symmetry is that there are three distinct diametrically opposite points and two distinct equilateral triangles that can be constructed using the six vertices of a regular hexagon. And more generally, given the geometrical fact that for natural numbers $m$ and $n$, we can find $m$ diameters in $2m$ sided regular polygons and $n$ equilateral triangles using the vertices of a $3n$ sided regular polygon only, it follows that in systems where anisotropic flows of order $6k$ dominate, for a natural number $k$, we can expect that the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ would have non-vanishing magnitudes with varying centrality and spherocity. The converse will also hold true, i.e., if the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ are comparable and non-vanishing, then $v_{6k}$ is dominant in the system for some natural number $k$, since there exist no other regular polygons, in which, few vertices can be chosen ensuring that the straight line connecting them is either a diameter or that they form an equilateral triangle. For the case when $k=1$, the above discussion holds true for hexagonal symmetry in the final state transverse momentum distribution. As can be seen, however, in Fig. 3(a), for the most central event classes, there is an overlap between the values of the SPCs in the different event shape classes. This is indicative of the fact that this correlator takes on its values due to event-by-event fluctuations in the Pb-Pb collision system rather than due to any dominant sixth-order anisotropy. More particularly, comparing Fig. 5(a) and Fig. 3(b), we can observe that although the SPC $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ has a qualitatively similar centrality dependence as that of $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ for the mid-central classes, it has different behaviour in the most central classes. From the inference previously drawn from $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, it can be understood that the values $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ takes is not due to an underlying anisotropy with hexagonal symmetry, but that particle emission in Pb-Pb collisions is largely dominated by $\psi_{2}$, as all other anisotropic flow terms only have a weak contribution.
However, if second and third-order anisotropies dominate in the system, then no such trivial link can be deduced. More specifically, if $v_{2}$ or $v_{3}$ dominate in comparison to $v_{6}$, then there are a few $6^{\rm th}$ order symmetry planes close to which the $2^{\rm nd}$ and $3^{\rm rd}$ order symmetry planes can lie in the transverse momentum space, which is the case for of Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. Although from the above discussion, the magnitude of the anisotropic flow coefficients seems to be playing an important role in the value that the SPCs take on, it is important to note that the anisotropic flow coefficient magnitudes have only been used to represent the symmetry of the anisotropic flow, and their contribution to SPCs has been removed by suitably choosing the harmonics and their respective powers in Eq. (II.4).
Figure 4: Comparison of SPC $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ between that calculated from events in which $v_{6}$ dominates $v_{2}$ and $v_{3}$, and all other events containing at least five charged hadrons with $p_{\rm T}>0.15$ GeV/c.
To make it clearer, a comparison has been shown in Fig. 4 between the SPC $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$ which has been obtained using all events in a centrality class, to that obtained using events with higher values of $v_{6}$. We can observe no positive shift in the values, owing to the geometrical fact that the particles are not ejected with transverse momenta distributed in a hexagonal symmetry. Gauging further into the mathematical construct of the SPCs, we find that a plethora of information about the transverse momentum distribution can be derived using SPCs of planes of lower harmonic orders. Following from the discussion at the beginning of Section III, if hexagonal symmetry is dominant in the system, then $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{3},\frac{2\pi}{3}\}$. Thus, all possible values for $6(\delta_{6}-\delta_{2})$ are of the form $2m\pi$ for some natural number $m$. Similarly, $|\delta_{6}-\delta_{3}|\in\{0,\frac{\pi}{3}\}$, and again, $6(\delta_{6}-\delta_{3})$ will always be of the form $2m^{\prime}\pi$ for some natural number $m^{\prime}$. However, when considering a system where, let us say, a $12^{\rm th}$ order symmetry is dominant, there, $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{2\pi}{3}\,\frac{5\pi}{6}\}$, so that $6(\delta_{6}-\delta_{2})\in\{0,\pi,2\pi,3\pi,4\pi,5\pi\}$. So, when these pass as arguments into the cosine function in the RHS of Eq. (II.4), some terms in the numerator would be positive while some negative, so that the average in the numerator will tend towards zero, implying in turn that the symmetry planes are decorrelated. This, however, is not the case, and in order to account for the correlation between symmetry planes of harmonic orders $n=2$ and $n=3$, in a system where a $12^{\rm th}$ order symmetry dominates, we must study the SPC $\langle\cos[12(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$. If the anisotropy in the system has very high harmonic orders of symmetry planes, then that would require us to study the SPC $\lim_{k\to\infty}\langle\cos[6k(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which when we try to find out, starting from small natural numbers $k$, would give us vanishingly low SPC values. However, only the small $k$ values are of any physical relevance since energy occurs clustered in nuclei. It is also important to keep in mind that SPCs would invariably yield low values for an isotropic distribution of the transverse momentum vectors. In this way, a careful choice of SPCs would allow us to gauge the higher-order anisotropies of the system. It is also necessary to emphasize, however, that the errors that would be associated with finding out the dominance of higher-order symmetries in the anisotropy of the distributions are not much lesser than what we would expect by computing the anisotropic flow coefficients themselves. This is because the SPC would require a larger number of particles to be correlated as well. The usefulness of using the SPCs, however, is that we can quickly reach a conclusion about the symmetry of the dominant anisotropic flow in the system by computing each anisotropic flow coefficient and comparing them individually.
Figure 5: Centrality dependence of SPCs (upper panel) for $n=2,4,6$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
Figure 6: Comparison between the RMS values of anisotropic flow coefficients of harmonic orders $n=2,4$ and 6 in different centrality classes in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV with events generated using AMPT.
III.2.2 SPCs of harmonic orders $n=2,4,6$
The fourth-order anisotropy would make the transverse momentum space geometry of the event appear like a square. Also, for a square, we have two distinct diagonals. These diagonals correspond to a second-order symmetry, that is, one corresponding to elliptic flow. Hence, a dominant presence of $v_{4}$ would imply that the symmetry planes $\psi_{4}$ and $\psi_{2}$ are strongly correlated. Hence, with decreasing centrality, the correlation between $\psi_{4}$ and $\psi_{2}$ is expected to increase, as is shown in Figs. 5(a) and (c). For the corresponding comparison, a plot of $v_{2}$, $v_{4}$ and $v_{6}$ is shown in Fig. 6. Furthermore, when $6^{\rm th}$ order anisotropy is dominant, it has already been described in Section III.2.1 that the correlation between $\psi_{6}$ and $\psi_{2}$ can be expected to be strong. The presence of strong hexagonal symmetry also brings with it the presence of an approximate quadrangular symmetry and, in turn, an elliptical symmetry of the anisotropy of the transverse momentum distribution of the final charged hadrons. Thus, in the presence of high $v_{4}$ and $v_{6}$, it can be expected that the symmetry planes of orders $n=2,4,6$ are well correlated, as can be seen from the comparison of Fig. 6 and Figs. 5(a) and (c).
Figure 7: Centrality dependence of SPCs (upper panel) among $n=2,3,5$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
III.2.3 SPCs of harmonic orders $n=2,3,5$
Whereas a pentagonal symmetry favors an approximate triangular symmetry (isosceles, but not equilateral triangles that can possibly be constructed using the vertices of a regular pentagon), the presence of elliptical symmetry is expected to be largely independent of either of the other two types of symmetries, when higher order anisotropies (e.g. $v_{30}$) are not dominant. Furthermore, as the isosceles triangles constructed out of a regular pentagon are not equilateral, we cannot expect any strong correlation between $\psi_{3}$ and $\psi_{5}$ as well. A strong SPC between harmonics $n=2,3,5$, even when different linear combinations of the symmetry plane angles are considered, would be indicative of the presence of a strong $15^{\rm th}$ order symmetry in the anisotropy of the transverse momentum distribution. As can be seen from the plots in Figs. 7(a) and (c), for two different SPCs for the same choice of harmonics, $n=2,3,5$, we see that the correlation is different, and in addition, the SPC $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ has values very close to zero.
III.3 Spherocity dependence of SPCs and PPCs
In Figs. 35, and 7, for most of the SPCs, regardless of their centrality, it is observed that the correlators for low-$S_{0}$ (pencil-like) case have a higher magnitude of the correlation compared to $S_{0}$-int events. This is also trivially expected, because the azimuthal angles that are considered in the averaging of the multiplets in the RHS of Eq. (I) are closer to each other for low-$S_{0}$ case, and the combination $\{n_{i}\}_{i=1}^{l}$ where $\sum_{i=1}^{l}n_{i}=0$, ensures that the argument that goes in as phase into the complex exponential in the RHS of Eq. (I) is closer to zero, leading to real parts that are closer to 1. Hence, all the numerator terms during event averaging using only low-$S_{0}$ events in Eq. (II.4) will have values that are more positive in comparison to those where the event averaging is done considering all shapes of events. Thus, if an SPC computed using events of all shapes is positive in a centrality class, it would mean that the same SPC would have a higher magnitude when computed using low-$S_{0}$ events.
On the other hand, when SPCs for the high-$S_{0}$ (isotropic) events, we have that the azimuthal angles are uniformly distributed, and hence the magnitude of SPCs evaluated using high-$S_{0}$ events would be smaller than that of $S_{0}$-int case. Apart from this triviality in terms of magnitude, no particular behaviour in terms of whether an SPC evaluated from isotropic events in a particular centrality class - whether it would be more positive or negative in comparison to the SPC obtained by considering events of all shapes - can be deduced from the mathematical construct of the correlators.
A deviation from this trivial behaviour would be indicative of interesting underlying phenomena that lead to the geometrical distribution of particles in the transverse momentum space. The SPCs that deviate from their trivially expected values when computed using only low-$S_{0}$ or high-$S_{0}$ events are $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{3})\rangle_{\rm GE}$, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$.
The participant planes are determined by the initial geometry of the colliding system. These are independent of the later stages of the evolution of the collision system. Hence, the participant planes are expected to be independent of spherocity of an event, which largely takes its value due to the later stages of evolution. Consequently, any correlation between the participant planes is also expected to be independent of the spherocity classification of an event. This is exactly what is observed in all the PPCs, which are shown in Figs. 5(d)-(f), Figs. 3(d)-(f), Figs. 7(d)-(f). The values of the PPCs match within errors for the averages that have been computed using the different kinds of event shapes. Their dependence on centrality is similar to that of the corresponding SPCs in Figs. 3(a)-(c), Figs. 5(a)-(c), Figs. 7(a)-(c).
For PPCs that grow with decreasing centrality, shown in Figs. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e), a striking difference can be seen between their spherocity dependence. A noticeable feature is that the SPCs mimic qualitatively the centrality dependence of the PPCs for all correlations, at least up to mid-central events (40-50% most central). Also, there are some other SPCs, e.g, $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{6})\rangle_{\rm GE}$ for which, the SPCs that are computed using low-$S_{0}$ events in the respective centrality class, mimic the qualitative centrality dependence of their corresponding PPCs, even for peripheral events.
The general behaviour of increase of these PPCs as well as SPCs, when we move towards mid-central event classes, is that higher-order anisotropies start playing a dominant role. Hence, the lower-order symmetry planes get correlated to each other non-trivially. For most central events, that are isotropic in transverse momentum space, there need not be any correlation between the different harmonic orders of the symmetry planes, i.e., the orientations of the symmetry planes need not be correlated among themselves at all. This is exactly what is observed in both the SPCs and the PPCs, as shown in Fig. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e)
Another comparison of SPCs to their respective PPCs can be made in Fig. 5(d) and Fig. 3(e), for PPCs that decrease as we move towards peripheral collisions. The reason for which $\langle\cos[6(\Phi_{3}-\Phi_{6})]\rangle_{\rm GE}$ decreases with centrality is due to its geometrical detail, as has been pointed out in Section III.2.1. For the PPC between the harmonic orders $n=2,3$ and 4, we have the fact that in most central collisions, the contribution of $v_{2}$ to anisotropic flow is less in comparison to that in peripheral events. As $\psi_{2}$ and $\psi_{4}$ together have an appreciable correlation with $\psi_{6}$ (Figs. 5(d) and (f)) and $\psi_{3}$ too has an appreciable correlation with $\psi_{6}$ (Fig. 3(d)), we observe a similar trend with centrality, for the simplest possible combination that has been constructed for the PPCs corresponding to $n=2,3$ and 4.
In addition to the SPCs and PPCs discussed above, Fig. 3(f) and Fig. 7(f) show the comparison of SPCs to the corresponding PPCs which are closer to zero within error bars across all centrality classes. These correlations have very small magnitudes and vary randomly about zero. This trend applies to their respective SPCs as well. However, in these cases, it is observed that the SPCs have a higher magnitude in comparison to the PPCs, even though their variation is random about the $x-$axis. In addition, no particular behaviour in terms of event shape can be observed for the PPCs and SPCs shown in Fig. 3(c),(f) and Fig. 7(c),(f). This is evidence that the correlator is actually close to zero irrespective of the event shape chosen, and hence the SPCs obtained using averaging over pencil-like or isotropic events show no specific centrality behaviour. This behaviour of the SPCs has been reported in the ALICE experiment [23]. In all the various comparisons that were made in this section, between PPCs and their corresponding SPCs, we observe a clear qualitative match of their centrality dependence. This shows, thus, that the SPCs are a good probe for PPCs.
IV Summary and Conclusion
In summary, we study centrality and transverse spherocity dependence of symmetry plane correlations in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. For the first time, this paper presents a study of SPCs using AMPT and event shape dependence of SPCs. The Gaussian estimator of SPCs has been observed to be strongly sensitive to event shape, and their variation with spherocity can be used to study them in greater detail in different centrality classes of interest, as per the physics process being addressed. When SPCs are computed using isotropic events, they have been observed to have smaller magnitude and random variation about zero, showing no particular centrality dependence, except for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$, where, it can be distinctly observed that the SPC obtained using isotropic events only takes on negative values, irrespective of the centrality. Most of the SPCs, except for those that are very close to zero, when computed for the low-$S_{0}$ events, are all larger than the $S_{0}$-int case. Thus, these will be particularly useful in examining if a certain SPC is actually zero due to the underlying geometry of the system or if it is zero simply due to it being computed from the chosen sample of events with all kinds of event shapes. Symmetry planes of even orders are likely to have greater correlation, as has been seen for the SPCs of $n=2,4,6$. All the SPCs, especially for the low-$S_{0}$ case, qualitatively mimic the centrality dependence of their corresponding PPCs. The PPCs do not depend strongly on spherocity, as they are not affected by the several processes that ultimately lead spherocity to take on the value that it does for an event. The qualitative estimation of PPCs by the SPCs will be an important tool to determine the relevant final state effect of nuclear structure at high-energy heavy-ion collisions, such as in 16O-16O collisions at the RHIC and LHC.
Acknowledgment
S. T. acknowledges the KVPY SB scholarship. S. P. acknowledges the University Grants Commission (UGC), Government of India, for doctoral fellowship. R. S. sincerely acknowledges the DAE-DST, Government of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). The authors gratefully acknowledge the computing facility provided by the Grid Computing Facility, VECC, Kolkata, India.